OpenStudy (anonymous):

College Algebra: I'm doing 2X2 Linear Systems and I'm not so well with fractions. Can someone solve this problem step by step for me? I'm reviewing for my chapter test coming up next week. I need to solve the system and determine whether it's independent, dependent, or inconsistent. 1/2x-1/3y=12 1/4x-1/2y=1 I'll rewrite the problem again below.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x-\frac{ 1 }{ 3 }y=12\] \[\frac{ 1 }{ 4 }x-\frac{ 1 }{ 2 }y=1\]

OpenStudy (triciaal):

to get rid of the fraction multiply each term by the LCM (lowest common multiple) of the denominator

OpenStudy (triciaal):

for equation 1 1/2 and 1/3 the lcm is 2*3 = 6 so multiply by 6 to get 3 x - 2 y = 72 now do equation 2 before trying to solve simultaneously

OpenStudy (triciaal):

what do you have for equation 2?

OpenStudy (anonymous):

I'm sorry, but I'm entirely lost. That doesn't seem like anything you're suppose to do. I need to try to combine the equations and solve.

OpenStudy (anonymous):

Then determine if it's independent, dependent, or inconsistent

myininaya (myininaya):

I think getting rid of the fractions like @triciaal as suggested would help you a ton especially since you said you don't like dealing with them. The cool thing about fractions in equations is you manipulate the equation so you see no fractions.

myininaya (myininaya):

Just whatever you do to one side an equation you can do the other.

OpenStudy (triciaal):

yes but it's easier to work with whole numbers so the first step is to get rid of the fraction

OpenStudy (triciaal):

to get rid of the fraction multiply each term in the equation by the lcm

OpenStudy (anonymous):

Okay, how did you get 3x-2y=72?

OpenStudy (anonymous):

I'm getting something else.

OpenStudy (triciaal):

I did equation 1 for you and now you need to do the same for equation 2

OpenStudy (anonymous):

I can'y do equation 2 if I don't know how you got equation 1.

OpenStudy (anonymous):

*can't

OpenStudy (triciaal):

repost for equation 1 1/2 and 1/3 the lcm is 2*3 = 6 so multiply by 6 to get 3 x - 2 y = 72 now do equation 2 before trying to solve simultaneously

myininaya (myininaya):

\[\frac{ 1 }{ 2 }x-\frac{ 1 }{ 3 }y=12 \text{ so I see 2 and 3 on bottom } \\ \text{ so I will multiply both sides by 2*3 } \\ 2 \cdot 3 \frac{1}{2}x-2 \cdot 3 \frac{1}{3}y=2 \cdot 3 (12) \\\cancel{2} \cdot 3 \frac{1}{\cancel{2}}x-2 \cdot \cancel{3} \frac{1}{\cancel{3}}y=2 \cdot 3 (12) \] since 2/2=1 and 3/3=1 so you have for the first equation can be written as 3x-2y=72

OpenStudy (anonymous):

Okay, now I see how you did that. Thank you. Now I can do equation 2.

myininaya (myininaya):

yep you try

OpenStudy (anonymous):

2x-4y=8

OpenStudy (anonymous):

Alight, so now I got the equations 3x-2y=72 and 2x-4y=8. What do I do next?

OpenStudy (triciaal):

for equation 2 do you see that you can divide by 2 it will be the same as x-2y = 4

OpenStudy (anonymous):

Okay, I see.

OpenStudy (triciaal):

there is nothing wrong with what you have you just didn't use the lowest common denominator for 2 and 4 which would be 4

OpenStudy (anonymous):

Oh, so you move the -2y on the other side. So it would be x=2y+4 right

OpenStudy (triciaal):

to solve by substitution yes

OpenStudy (anonymous):

That's what I'm doing at the moment. :) Awesome!

OpenStudy (triciaal):

there are 4 ways to solve simultaneous equations. substitution is just one of them

OpenStudy (anonymous):

3(2y+4)-2y=72 is what I got. Then I got 6y+12-2y=72. Then combine like-terms and it will be 4y+12=72. Subtract on both sides and it's 4y=60. Divide on both sides and it's y=15

OpenStudy (triciaal):

now you substitute for x in equation 1 and solve for y use this value in the expression for x to get the value for x

OpenStudy (triciaal):